# Laser Interactions with Gas Jets: EMP Emission and Nozzle Damage

### v1

@@ I Introduction
A second important concern of the laser-gas interaction is damage to the gas jet nozzle. Many of the most exciting applications of laser-gas research rely on high shot rates and a reproducible gas density profile to generate bright, high-fluence sources of energetic ions [11, 12, 13, 14, 15, 16], electrons[17, 18] or x-rays [19, 20]. Repeated melting of the gas nozzle therefore represents a serious and expensive hindrance to this research. On energetic systems like the Vulcan Petawatt laser [21, 22] at the Rutherford Appleton Laboratory, a single shot on a dense gas target is sufficient to destroy the gas nozzle completely (see Fig. 1). For systems operating at lower energy, lower gas density and shorter pulse duration, nozzle damage is more progressive but still leads to significant smoothing of the gas density profile [23, 24], degradation of the laser interaction and reduced data reproducibility. Understanding the nozzle damage mechanism is therefore important for future laser-gas applications.
@@
To date, no theoretical framework has been proposed for the emission of EMP from gas jets and this makes it difficult to estimate the severity of its impact on new laser systems. In Sec. II and Sec. III.1-III.3, we present a model of laser-plasma expansion in a gas, where ionisation is mainly caused by ions streaming from the laser focus (anode) to the nozzle (cathode). In Sec. III.4, however, we show that electrical breakdown induced by a strong plasma-nozzle potential can fully ionise the gas on shorter timescales of 10-20 ps. EMP is then emitted similarly to solid target interactions, where a discharge current propagates to ground (the chamber) along a cm-scale antenna (the nozzle or target support), emitting radiation at gigahertz frequencies. According to our model, the EMP amplitude for Vulcan Petawatt and VEGA-style interactions can reach $\sim 100$ $\mu$T or $\sim 100$ kV/m at a distance $\sim 1$ m from the target. The theoretical model is supported in Sec. VI by the results from two experiments on the VEGA-3 laser, which show reasonable agreement with the discharge time, magnetic field spatial profile and accumulated target charge.
Using the plasma expansion model in Sec. V, we consider how the nozzle may be damaged. Two mechanisms are considered: (i) a kA-level discharge current and (ii) ion collisional heating. We show that damage to gas jet nozzles is more likely caused by the impact of plasma ions than Ohmic heating of the nozzle surface. Finally, in Sec. VII, we present ideas for how the EMP emission model can be reliably benchmarked with simulations and dedicated experiments. The practical impact of this work is broad: allowing scientists to reduce damage to expensive gas jet nozzles and minimize the electrical disruption of equipment. It likewise represents a rich seam of more fundamental research, connecting the physics of laser-target charging, ionisation, ion acceleration, high-voltage breakdown and antenna emission.
@@ II Expansion of a Laser-Plasma in a Gas Jet
We consider an expanding plasma created as a laser pulse propagates through a high-density gas. The laser pulse ionises the gas and creates a plasma channel. The plasma electrons are heated by the laser radiation to relativistic energies, and some of them escape the channel and leave it positively charged. This charged plasma cylinder then expands, ionising the gas until it makes contact with the conducting tip of the gas nozzle. Once contact has been made, the plasma is connected directly to the ground via the nozzle. A discharge is triggered and EMP radiation is emitted. Nozzle damage is determined variously by the amount of energy stored in the plasma, the strength of the discharge and the nozzle material.
@@
The Rayleigh length of a diffraction-limited Gaussian beam in our example is $Z_{R}=\pi w_{L}^{2}/\lambda_{L}\simeq 44\,\mu$m. Since the gas jet width is usually larger than $Z_{R}$, the laser produces a channel a few times longer than $Z_{R}$ that can be viewed as a $l_{p}\simeq 0.5$ mm-long plasma cylinder created at a typical height of $h\simeq 0.5$ mm above the nozzle. We take the initial radius of the plasma channel to be a factor of a few times larger than the laser focal radius, $r_{p}=20\,\mu$m, to reflect the average size of the channel as the laser deposits energy over the full length of its path through the gas. The laser absorption in the gas jet is $\eta_{L}\sim 90$%, with 40% transferred to hot electrons. This is consistent both with Particle-in-Cell (PIC) simulations[25, 26] and experimental interferograms [27], which show that a high-power laser pulse deposits a large fraction of its energy in the gas jet and leaves behind it a plasma channel several tens of microns wide.
@@
Table 1: Representative experimental parameters - as described in the text and Refs. [25, 23] - and estimates of plasma characteristics according to the model presented in Sec. II. The plasma radius $r_{p}$ and the laser energy fraction converted to hot electrons $\eta_{\rm las\rightarrow e}$ are estimated from dedicated PIC simulations. Values are given to 2 significant figures.
@@
Figure 2: Schematic of the nozzle, gas and the cylindrical plasma channel formed by the laser. Here, $r_{p}$ is the radius of the plasma cylinder, $h$ is the separation between the channel and the nozzle, and $\Phi_{p}$ is the electric potential. The laser is directed into the page, along the axis of the plasma. Red crosses indicate a region of positive charge inside the laser-generated channel. Three curved arrows show the geometry of the electric field, E.
@@ II.1 Electric Charging of the Plasma
Following the available observations and numerical simulations [28, 29, 30, 31], we assume that the plasma in the laser channel is fully ionised. The absorbed laser energy $\eta_{L}E_{L}$ is transferred to electrons with an energy distribution approximated by a two-temperature distribution:
@@
where $f_{\rm MJ}(\gamma,\Theta)$ is the Maxwell-Jüttner function
@@
depending on the dimensionless temperature $\Theta=T/m_{e}c^{2}$, the second order Bessel function $K$ and the electron relativistic factor $\gamma=1+\varepsilon/m_{e}c^{2}$. A fraction $\eta_{\rm las\to e}$ is transferred to a population of hot electrons with density $n_{h}$ and temperature $T_{h}$, while the remaining energy goes to electrons in the plasma bulk, with density $n_{e}$ and temperature $T_{e}$.
The bulk electron density is defined by the gas density ($n_{e}=n_{\rm at}$ for hydrogen) and the bulk temperature is defined by the plasma volume and deposited energy via:
$C_{e}(T_{e})n_{e}T_{e}=(1-\eta_{\rm las\to e})\eta_{L}E_{L}/V_{p},$ (2)
@@
To estimate the hot electron density, we prescribe a hot electron temperature of $T_{h}=5.0$ MeV following the ponderomotive scaling[32]. The hot electron density can then be defined knowing the fraction of energy transferred to hot electrons $\eta_{\rm las\to e}$:
$C_{e}(T_{h})n_{h}T_{h}=\eta_{\rm las\to e}\eta_{L}E_{L}/V_{p}.$
On Vulcan, the hot electron fraction is $n_{h}/n_{e}\simeq 2.5$% and their total energy is 29 J.
The electric potential of the plasma channel is determined by how many hot electrons can overcome the plasma potential barrier. The electric potential $\Phi_{p}$ at the surface of the plasma cylinder with respect to the grounded nozzle can be estimated as
@@
where $Q$ is the charge of the plasma and $C=2\pi l_{p}\epsilon_{0}/\ln(h/r_{p})$ is the capacitance. Only electrons with energies $\varepsilon_{e}>e\Phi_{p}$ can escape plasma to the ground. Assuming, as mentioned above, that hot electrons have an exponential distribution in energy with a temperature $T_{h}$, the fraction of escaped hot electrons is [33]:
@@
where $\psi=e\Phi_{p}/m_{e}c^{2}$ is the dimensionless potential. The plasma charge is then $Q=e\delta n_{h}V_{p}$ and, substituting this expression into Eq. (3), we have the following equation for the dimensionless potential $\psi$:
@@
where $\omega_{ph}=(e^{2}n_{h}/m_{e}\epsilon_{0})^{1/2}$ is the hot electron plasma frequency. Solution of this equation for $h/r_{p}=250$ and $r_{p}\omega_{ph}/c=18$ gives $\Phi_{p}=24.8$ MV. Knowing the plasma potential and the plasma capacitance $C\simeq 0.0086$ pF, we find the charge ejected from the plasma column, $Q=C\Phi_{p}\simeq 234$ nC. This corresponds to a fraction of escaped electrons $\delta n_{h}/n_{e}\approx 2.8\times 10^{-3}$ and is consistent with the experiment [34]. Since the time required for a hot electron to cross the channel is relatively short ($\sim 2r_{p}/c\sim 0.13$ ps), the time of plasma charging is determined by the time of flight of the laser pulse through the gas, $\sim 3$ ps.
@@ II.2 Ion Acceleration in Expanding Plasma
After the end of the laser pulse, there is no more energy supply and the plasma expands adiabatically under the electron thermal pressure. For the plasma parameters given above, the pressure $p_{e}=n_{e}T_{e}\simeq 330$ Mbar is much higher than the ambient gas pressure, so the plasma expands freely with electrons transferring their energy to ions. The plasma charge is conserved during the expansion phase and the potential decreases logarithmically as the plasma radius increases (see Eq. (3)). The plasma expansion can be described in some special cases by a self-similar rarefaction wave model [35, 36]. More detailed analysis is performed for a spherical plasma expansion in Ref. [37]. The ion energy distribution depends on the density profile and on the ratio of the plasma radius to the Debye length. For our crude estimates, it is sufficient to estimate the average ion kinetic energy by dividing the deposited laser energy by the number of ions in the plasma. For a hydrogen plasma, it is 1.5 times the bulk electron temperature $\varepsilon_{i}\simeq 0.4$ MeV. This corresponds to an ion velocity of $v_{i}\simeq 9\,\mu$m/ps, and an ion expansion time to the nozzle $t_{\rm exp}\simeq 60$ ps. The total energy carried by these ions is of order the bulk electron energy - that is, about 40 J for the Vulcan experiment.
@@ III.1 X-ray Photoionisation of the Gas
The bremsstrahlung emission of electrons in the laser-created plasma covers a broad range of photon energies up to the electron kinetic energy. Assuming the electron energy distribution is characterized by an effective temperature $T_{e}$, the power of bremsstrahlung emission from a hydrogen plasma can be written as [38]:
@@
where the electron density is in cm-3 and the electron temperature is in eV. For an electron density of $n_{e}=8\times 10^{20}$ cm-3 in fully ionised plasma and a temperature $T_{e}\simeq 250$ keV, we have $P_{\text{Brem}}=6\times 10^{12}$ W cm-3. Accounting for the plasma volume of $V_{p}\sim 6\times 10^{-7}$ cm3 and lifetime $t_{\rm exp}\sim 60$ ps, the total radiated energy is about $\mathcal{E}_{\rm Brem}=P_{\rm Brem}V_{p}t_{\rm exp}\simeq 0.2$ mJ. The spectrum of emission is flat for photons with energy $\varepsilon_{x}\lesssim T_{e}$ and decreases exponentially with a temperature $T_{\rm rad}\simeq T_{e}$ for higher energies.
@@ III.4 Gas Ionisation by Electrical Breakdown
When hot electrons are ejected from the laser plasma, it becomes positively charged with respect to the nozzle. The electric field between the plasma and nozzle may then break down the gas and propagate a discharge. The discharge regime is determined by the product of the gas pressure and electrode separation $p_{\rm gas}h$. Under Vulcan-like conditions, with $p_{\rm gas}h\gtrsim 1$ bar cm, avalanche ionisation will occur if the potential is above the breakdown value of $20-30$ kV. This is much smaller than the plasma potential of $\sim 25$ MV estimated in Sec. II.
In the avalanche process, the electron density increases exponentially in time from the seed level $n_{e0}/n_{\rm at}\sim 10^{-4}-10^{-3}$ created by fast electrons or ions, $n_{e}(t)=n_{e0}\,\exp(t/\tau_{\rm ion})$, where $\tau_{\rm ion}$ is the characteristic ionisation time. A high electrostatic potential means the ionisation time is very short, $\tau_{\rm ion}\sim 1$ ps. However, the plasma potential decreases with time for two reasons: first, it decreases logarithmically with time because of plasma expansion according to Eq. (3) and second, because the energy deposited during the ionisation process is extracted from the electrostatic energy. The decrease in electrostatic energy is related to the ionisation loss, $\simeq U_{\rm ion}n_{e}V_{\rm gas}$, and the discharge stops when $\mathcal{E}_{es}\sim U_{\rm ion}V_{\rm gas}n_{e}$. This relation defines the maximum ionisation that can be achieved by avalanche breakdown, $n_{e,\max}\sim 4\times 10^{21}$ cm-3, which is larger than the atomic gas density.
In conclusion, avalanche breakdown produces a higher level of ionisation than fast electrons or fast ions on a $10-20$ ps timescale, and it can fully ionise the gas under the conditions of the Vulcan experiment before the expanding plasma reaches the nozzle.

### v2

@@ I Introduction
A second important concern of the laser-gas interaction is damage to the gas jet nozzle. Many of the most exciting applications of laser-gas research rely on high shot rates and a reproducible gas density profile to generate bright, high-fluence sources of energetic ions [11, 12, 13, 14, 15, 16], electrons[17, 18] or x-rays [19, 20]. Repeated melting of the gas nozzle therefore represents a serious and expensive hindrance to this research. On energetic systems like the Vulcan Petawatt laser [21, 22] at the Rutherford Appleton Laboratory, a single shot on a dense gas target is sufficient to destroy the gas nozzle completely (see Fig. 1). For systems operating at lower energy, lower gas density and shorter pulse duration, nozzle damage is more progressive but still leads to significant smoothing of the gas density profile [23, 24, 25], degradation of the laser interaction and reduced data reproducibility. Understanding the nozzle damage mechanism is therefore important for future laser-gas applications.
@@
To date, no theoretical framework has been proposed for the emission of EMP from gas jets and this makes it difficult to estimate the severity of its impact on new laser systems. In Sec. II and Sec. III.1-III.3, we present a model of laser-plasma expansion in a gas, where ionisation is mainly caused by ions streaming from the laser focus (anode) to the nozzle (cathode). In Sec. III.4, however, we show that electrical breakdown induced by a strong plasma-nozzle potential can fully ionise the gas on shorter timescales of 10-20 ps. EMP is then emitted similarly to solid-target interactions, where a discharge current propagates to ground (the chamber) along a cm-scale antenna (the nozzle or target support), emitting radiation at gigahertz frequencies. According to our model, the EMP amplitude for Vulcan Petawatt and VEGA-style interactions can reach $\sim 100$ $\mu$T or $\sim 100$ kV/m at a distance $\sim 1$ m from the target. The theoretical model is supported in Sec. VI by the results from two experiments on the VEGA-3 laser, which show reasonable agreement with the discharge time, magnetic field spatial profile and accumulated target charge.
In Sec. V, we consider how the nozzle may be damaged using the plasma expansion model. Two mechanisms are considered: (i) a kA-level discharge current and (ii) ion collisional heating. We show that damage to gas jet nozzles is more likely caused by the impact of plasma ions than Ohmic heating of the nozzle surface. Finally, in Sec. VII, we present ideas for how the EMP emission model can be reliably benchmarked with simulations and dedicated experiments. The practical impact of this work is broad: allowing scientists to reduce damage to expensive gas jet nozzles and minimize the electrical disruption of equipment. It likewise represents a rich seam of more fundamental research, connecting the physics of laser-target charging, ionisation, ion acceleration, high-voltage breakdown and antenna emission.
@@ II Expansion of a Laser-Plasma in a Gas Jet
We consider an expanding plasma created as a laser pulse propagates through a high-density gas. The laser pulse ionises the gas and creates a plasma channel. Plasma electrons are heated by the laser radiation to relativistic energies, and some of them escape the channel and leave it positively charged. This charged plasma cylinder then expands into the surrounding gas or plasma, depending on the efficiency of the ionisation mechanisms discussed in Sec. III. If the plasma expands into gas, the gas is ionised until contact is made with the conducting tip of the gas nozzle. Once contact is made and the plasma is connected directly to ground, a discharge is triggered and EMP radiation is emitted. Nozzle damage is determined variously by the amount of energy stored in the plasma, the strength of the discharge and the nozzle material.
@@
The Rayleigh length of a diffraction-limited Gaussian beam in our example is $Z_{R}=\pi w_{L}^{2}/\lambda_{L}\simeq 44\,\mu$m. Since the gas jet width is usually larger than $Z_{R}$ and the beam is susceptible to relativistic self-guiding, the laser produces a channel a few times longer than $Z_{R}$ that can be viewed as a $l_{p}\simeq 0.5$ mm-long plasma cylinder created at a typical height of $h\simeq 0.5$ mm above the nozzle. We take the initial radius of the plasma channel to be a factor of a few times larger than the laser focal radius, $r_{p}=20\,\mu$m, to reflect the average size of the channel as the laser deposits energy over the full length of its path through the gas. We assume that laser absorption in the gas jet is $\eta_{L}\sim 90$%, with 40% transferred to hot electrons. This is consistent both with Particle-in-Cell (PIC) simulations[26, 27] and experimental interferograms [28], which show that a high-power laser pulse deposits a large fraction of its energy in the gas jet and leaves behind it a plasma channel several tens of microns wide.
@@
Table 1: Representative experimental parameters - as described in the text and Refs. [26, 24] - and estimates of plasma characteristics according to the model presented in Sec. II. The plasma radius $r_{p}$ and the laser energy fraction converted to hot electrons $\eta_{\rm las\to h}$ are estimated from dedicated PIC simulations.
@@
Figure 2: Schematic of the nozzle, gas and the cylindrical plasma channel formed by the laser. Here, $r_{p}$ is the radius of the plasma cylinder, $h$ is the separation between the channel and the nozzle, and $\Phi_{p}$ is the electric potential. The laser is directed into the plane of the page, along the axis of the plasma. Red crosses indicate a region of positive charge inside the laser-generated channel. Three curved arrows sketch the geometry of the electric field, E.
@@ II.1 Electric Charging of the Plasma
Following the available observations and numerical simulations [29, 30, 31, 32], we assume that the plasma in the laser channel is fully ionised. The absorbed laser energy $\eta_{L}E_{L}$ is transferred to electrons with an energy distribution approximated by a two-temperature relativistic distribution:
@@
where $f_{\rm MJ}(\gamma,\Theta)$ is the Maxwell-Jüttner function[33]
@@
depending on the dimensionless temperature $\Theta=T/(m_{e}c^{2})$, the modified Bessel function $K$ of the second kind and the electron relativistic factor $\gamma=1+\varepsilon/(m_{e}c^{2})$. A fraction $\eta_{\rm las\to h}$ of the absorbed laser energy is transferred to a population of hot electrons with density $n_{h}$ and temperature $T_{h}$, while the remaining energy goes to electrons in the plasma bulk, with density $n_{e}$ and temperature $T_{e}$. Here, we use subscripts “h” and “e” to refer to the hot and thermal electron populations respectively.
The bulk electron density is defined by the gas density ($n_{e}=Zn_{\rm at}$) and the bulk temperature is defined by the plasma volume and deposited energy via:
$C_{e}(T_{e})n_{e}T_{e}=(1-\eta_{\rm las\to h})\eta_{L}E_{L}/V_{p},$ (2)
@@
To estimate the hot electron density, we prescribe a hot electron temperature of $T_{h}=5.0$ MeV following the ponderomotive scaling[34], $(\sqrt{1+a_{0}^{2}}-1)m_{e}c^{2}$. The hot electron density can then be defined knowing the fraction of energy transferred to hot electrons $\eta_{\rm las\to h}$:
$C_{e}(T_{h})n_{h}T_{h}=\eta_{\rm las\to h}\eta_{L}E_{L}/V_{p}.$
Knowing the hot electron temperature and assuming that $\eta_{\rm las\to h}=0.4$ on Vulcan, this equation gives a hot electron fraction $n_{h}/n_{e}\simeq 2.5$% with total energy 29 J.
The electric potential of the plasma channel is determined by how many hot electrons can overcome the plasma potential barrier. Though electrons from the plasma bulk will also contribute to the ejected charge, their contribution is less than $10\%$ because a significant fraction of the available laser energy is transferred to hot electrons. The electric potential $\Phi_{p}$ at the surface of the plasma cylinder with respect to the grounded nozzle can be estimated as
@@
where $Q$ is the charge of the plasma and $C=2\pi l_{p}\epsilon_{0}/\ln(h/r_{p})$ is the capacitance. The capacitance follows from the potential of a charged cylinder, radius $r_{p}$, length $l_{p}$, situated a distance $h$ from an infinite conducting plane. Only electrons with energies $\varepsilon_{e}>e\Phi_{p}$ can escape plasma to the ground. Assuming, as mentioned above, that hot electrons have an exponential distribution in energy with a temperature $T_{h}$, the fraction of escaped hot electrons is [35]:
@@
where $\psi=e\Phi_{p}/(m_{e}c^{2})$ is the dimensionless potential. The plasma charge is then $Q=e\delta n_{h}V_{p}$ and, substituting this expression into Eq. (3), we have the following equation for the dimensionless potential $\psi$:
@@
where $\omega_{ph}=(e^{2}n_{h}/(m_{e}\epsilon_{0}))^{1/2}$ is the hot electron plasma frequency. The solution of this equation for typical parameters $h/r_{p}=250$ and $r_{p}\omega_{ph}/c=18$ gives $\Phi_{p}=24.8$ MV. Knowing the plasma potential and the plasma capacitance $C\simeq 0.0086$ pF, we find the charge ejected from the plasma column, $Q=C\Phi_{p}\simeq 234$ nC. This corresponds to a fraction of escaped electrons $\delta n_{h}/n_{e}\approx 2.8\times 10^{-3}$ and is consistent with the experiment [36]. Since the time required for a hot electron to cross the channel is relatively short ($\sim 2r_{p}/c=0.13$ ps), the time of plasma charging is determined by the propagation time of the laser pulse through the gas, $\sim 3$ ps.
@@ II.2 Ion Acceleration in Expanding Plasma
After the end of the laser pulse, there is no more energy supply and the plasma expands adiabatically under the electron thermal pressure. For the plasma parameters given above, the pressure $p_{e}=n_{e}T_{e}\simeq 330$ Mbar is much higher than the ambient gas pressure, so the plasma expands freely with electrons transferring their energy to ions. The plasma charge is conserved during the expansion phase and the potential decreases logarithmically as the plasma radius increases (see Eq. (5)). The plasma expansion can be described in some special cases by a self-similar rarefaction wave model [37, 38]. More detailed analysis is performed for a spherical plasma expansion in Ref. [39]. The ion energy distribution depends on the density profile and on the ratio of the plasma radius to the Debye length. For our purposes, it is sufficient to estimate the average ion kinetic energy by dividing the deposited laser energy by the number of ions in the plasma. For a hydrogen plasma, it is 1.5 times the bulk electron temperature $\varepsilon_{i}\simeq 0.4$ MeV. This corresponds to an ion velocity of $v_{i}\simeq 9\,\mu$m/ps, and an ion expansion time to the nozzle $t_{\rm exp}\simeq 60$ ps. The total energy carried by these ions is of order the bulk electron energy - that is, about $(1-\eta_{\rm las\to h})\eta_{L}E_{L}\sim 40$ J for the Vulcan experiment.
@@ III.1 X-ray Photoionisation of the Gas
The bremsstrahlung emission of electrons in the laser-created plasma covers a broad range of photon energies up to the electron kinetic energy. Assuming the electron energy distribution is characterized by an effective temperature $T_{e}$, the power per unit volume of bremsstrahlung emission from a hydrogen plasma can be written as [40]:
@@
where the electron density is in cm-3 and the electron temperature is in eV. The radiative contribution of bulk electrons dominates that of hot electrons because $n_{e}T_{e}\gtrsim n_{h}T_{h}$ under the present conditions. For an electron density of $n_{e}=8\times 10^{20}$ cm-3 in fully ionised plasma and a temperature $T_{e}\simeq 250$ keV, we have $P_{\text{Brem}}=6\times 10^{12}$ W cm-3. Accounting for the plasma volume of $V_{p}\sim 6\times 10^{-7}$ cm3 and lifetime $t_{\rm exp}\sim 60$ ps, the total radiated energy is about $\mathcal{E}_{\rm Brem}=P_{\rm Brem}V_{p}t_{\rm exp}\simeq 0.2$ mJ. The spectrum of emission is flat for photons with energy $\varepsilon_{x}\lesssim T_{e}$ and decreases exponentially with a temperature $T_{\rm rad}\simeq T_{e}$ for higher energies.
@@ III.4 Gas Ionisation by Electrical Breakdown
When hot electrons are ejected from the laser plasma, it becomes positively charged with respect to the nozzle. The electric field between the plasma and nozzle may then break down the gas and propagate a discharge. The discharge regime is determined by the product of the gas pressure and electrode separation $p_{\rm gas}h$. Under Vulcan-like conditions, with $p_{\rm gas}h\gtrsim 1$ bar cm, avalanche ionisation will occur if the potential is above the breakdown value of $20-30$ kV  [48]. This is much smaller than the plasma potential of $\sim 25$ MV estimated in Sec. II.
In the avalanche process, the electron density increases exponentially in time from the seed level $n_{e0}/n_{\rm at}\sim 10^{-4}-10^{-3}$ created by fast electrons or ions, $n_{e}(t)=n_{e0}\,\exp(\nu_{\rm ion}t)$, where $\nu_{\rm ion}$ is the characteristic ionization rate. According to Refs. [48, 49], the ionisation rate is $\nu_{\rm ion}\sim 10^{12}$ s-1 for Vulcan-like conditions and the gas can be fully ionised in a few picoseconds. However, the plasma potential decreases with time for two reasons: first, it decreases logarithmically with time because of plasma expansion according to Eq. (5) and second, because the energy deposited during the ionisation process is extracted from the electrostatic energy. The decrease in electrostatic energy is related to the ionisation loss, $\Delta\mathcal{E}\simeq U_{\rm ion}n_{e}V_{\rm gas}$, and the discharge stops when all the available electrostatic energy $\mathcal{E}_{es}$ has been exhausted. This relation defines the maximum ionisation that can be achieved by avalanche breakdown, $n_{e,\max}\sim 4\times 10^{21}$ cm-3, which is larger than the atomic gas density.
In conclusion, avalanche breakdown is initiated from the seed level of ionisation produced by fast electrons or fast ions on a $10-20$ ps timescale. It can fully ionise the gas under the conditions of the Vulcan experiment before the expanding plasma reaches the nozzle.